OpenStudy (anonymous):

Elizabeth had a book delivered to her home. The total cost included the price of the book, 7% sales tax on the price of the book, and a $4.50 shipping charge. Elizabeth paid a total of $31.25. What was the pre-tax price of the book? $

OpenStudy (faiqraees):

okay first calculate the value she had to pay only for the book and the tax

OpenStudy (anonymous):

26.75$

OpenStudy (faiqraees):

okay 26.75 is the value of the book and the 7% of the books cost. Agreed?

OpenStudy (anonymous):

yes.

OpenStudy (faiqraees):

We can form an equation \[x+ (0.07x)=26.75\]

OpenStudy (faiqraees):

7% =7/100 =0.07 Clear so far?

OpenStudy (anonymous):

yes.

OpenStudy (faiqraees):

solve for x Can you do it?

OpenStudy (anonymous):

I think it's 1.87.

OpenStudy (faiqraees):

no.. x+0.07x=26.75 1x+0.07x=26.75 1.07x=26.75 so whats the value of x now

OpenStudy (anonymous):

3.82?

OpenStudy (faiqraees):

divide 1.07 on both sides 1.07 26.75 -----x = ----- 1.07 1.07 x= 26.75 ------- 1.07

OpenStudy (anonymous):

were did you get a 1.07?

OpenStudy (faiqraees):

x+0.07x=26.75 1x+0.07x=26.75 1.07x=26.75

OpenStudy (anonymous):

ohhh okay.